Solve for x and do a formal check. 3(2x – 3) – 4(x – 2) = -(x – 5)
Answer by abdulimad2(2)

3(2x-3)-4(x-2) = -(x-5)
Start by distributing any variables into the expressions that have parentheses:
(tip: DO NOT FORGET TO DISTRIBUTE THE MINUS SIGN INTO THE EXPRESSION IF IT EXISTS)
=> 6x-9-4x+8 = -x+5
Combine like terms:
=> 6x-4x-9+8 = -x+5
=> 2x-1 = -x+5
Isolate the variable x on the left side of the equation, and the numbers on the right side of the equation:
=> 2x-1+x=5
=> 3x-1 = 5
=> 3x = 5+1
=> 3x = 6
=> x = 6/3
=> x = 2
After finding the value of x, we can substitute its value back into the ORIGINAL EQUATION to check.
=> 3(2(2)-3)-4(2-2) = -(2-5)
=> 3(4-3)-4(0) = -(-3)
=> 3(1) - 0 = 3
=> 3 = 3
Since we plugged 2 in place of x in the original equation and got that both sides are equal to each other (3=3), this means that our solution of x=2 is true. If we didn't get both sides to be equal to each other, then our answer would be wrong, and we'd have to go back to the equation and solve it again.